# The Mean of Multi-Object Trajectories

### v1

@@ Abstract
This paper introduces the concept of a mean for trajectories and multi-object trajectories–sets or multi-sets of trajectories–along with algorithms for computing them. Specifically, we use the Fréchet mean, and metrics based on the optimal sub-pattern assignment (OSPA) construct, to extend the notion of average from vectors to trajectories and multi-object trajectories. Further, we develop efficient algorithms to compute these means using greedy search and Gibbs sampling. Using distributed multi-object tracking as an application, we demonstrate that the Fréchet mean approach to multi-object trajectory consensus significantly outperforms state-of-the-art distributed multi-object tracking methods.
Index Terms:
mean, average, consensus, mean trajectory, mean multi-object trajectory, distributed multi-object fusion.
@@ I Introduction
Multi-object estimation involves determining the trajectories of an unknown, time-varying number of objects from noisy sensor measurements. Formally, a trajectory is a function that maps time indices from a discrete time window to elements in some state space, and a multi-object trajectory is a point pattern or a multi-set of trajectories [beard2020solution]. This general definition covers trajectories that are fragmented (i.e., objects can disappear and then reappear later), which is usually the case with the output of most tracking systems. Estimating the multi-object trajectory lies at the heart of many applications ranging from space exploration [persico2018cubesat], surveillance [liu2022multi, nguyen2019trackerbots], computer vision [ma2024track, zhang2021fairmot], to cell biology [mavska2023cell, nguyen2021tracking].
@@
While it is customary that the mean is simply the sum of the data points divided by the number of data points, this is not the case in multi-object estimation. When the data points are trajectories of different lengths, or multi-object trajectories, see e.g., Figure 1, the customary notion of average is not applicable because addition is not defined. The fundamental question then arises: what constitutes an “average” trajectory or an “average” multi-object trajectory? This question is crucial in multi-object estimation, yet, to the best of our knowledge, there is no existing work addressing the concept of an “average” trajectory or an “average” multi-object trajectory.
(a) 2D Fréchet Mean Trajectory
(b) 2D Fréchet Mean Multi-Object Trajectory
Figure 1: (a) Trajectories (color) and candidate means (thick black); marker sizes indicate times, with the same size representing the same time, and sizes decrease toward the ends of trajectories (b) Multi-object trajectories (color) and candidate means (thick black); each multi-object trajectory is given the same marker type with member trajectories distinguished by different colors.
@@
Additionally, we demonstrate the capability of the proposed means in achieving consensus, using networked multi-object tracking as a specific area of application. Due to the proliferation of low-cost sensors that enables the deployment of large networks capable of monitoring large spatial regions, networked multi-object tracking is an emerging trend [iyer2022wind]. This approach leverages advances in communication and sensing technologies to improve accuracy by fusing information from multiple nodes in a flexible and reliable manner, resilient to node failures [luo2006distributed]. Given the local multi-object trajectory descriptions (e.g., estimates, probability distributions) at each node of the network, fusion is used to determine a multi-object trajectory description that best agrees with the local nodes, which can be posed as a consensus on the local multi-object trajectories. The proposed consensus can fuse node outputs regardless of the types of local trackers, and shows significant performance improvements over current state-of-the-art techniques.
@@ II-A Distributed Multi-Object Estimation
An alternative to multi-scan multi-object density fusion is track-to-track (T2T) fusion [chang1997optimal]. This is an intuitive technique for fusing individual tracks together by minimizing some prescribed cost functions. Additional information on the state can be augmented into the cost function [osbome2011track]. A popular cost function is based on the association probability of the sets of estimated objects with the physical objects [bar1981track, ogle2019alternative]. Alternatively, some authors use cost functions based on the statistical distance/divergence between the distributions describing the states of the tracks and their associated confidence [tang2021multi, battistelli2014kullback, tang2018information], or the OSPA distance between the tracks [nguyen2021distributed]. T2T assignment can be cast as a multi-dimensional assignment problem, which is NP-hard for more than two dimensions. Nevertheless, sub-optimal solutions can be obtained using efficient clustering techniques, including hierarchical clustering [bereson2006efficient], DBSCAN [he2018multi], or stochastic optimization [wolf2023track]. In T2T fusion, clutter objects are typically not considered within the optimization process but are instead discarded based on existence probability [muvsicki2015track], voting logic [zhang1997efficient], or assignment confidence [stankovic2021adaptive]. Determining an effective cost function for T2T fusion is still an active research topic [wang2022consistent, kaplan2008assignment, bar1986effect].
Numerically, multi-object trajectory fusion is more efficient with T2T fusion since only the individual states or their distributions are needed, rather than the entire multi-object densities. In [bar2004multisensor], the multi-scan assignment problem was solved using Lagrangian relaxation for a small number of sensors. A fast algorithm for multi-scan T2T assignment was proposed in [tokta2018fast], but it is limited to two sensors. The OSPA-based multi-object trajectory fusion approach proposed in [nguyen2021distributed] has been shown to be effective for a large number of sensors. Further, as a peer-to-peer fusion algorithm, it is more robust to network disruption. However, it is less accurate than global fusion algorithms, wherein all information are fused at once.
@@ II-B Metrics for Point Patterns and Trajectories
Since a multi-object state is a point pattern, metrics for point patterns are natural for measuring the distance between multi-object states. While the Hausdorff metric is an obvious point pattern metric, it is insensitive to cardinality differences [hoffman2004multitarget]. To address this, the optimal mass transfer metric was proposed in [hoffman2004multitarget], based on the Wasserstein construction. However, this metric suffers from a physical inconsistency if the cardinalities of the two multi-object states are not the same [hoffman2004multitarget]. A widely used multi-object metric with meaningful physical interpretations, capturing both the state and cardinality errors, is the OSPA metric [schuhmacher2008consistent]. The OSPA construction also admits other metrics such as the cardinalized optimal linear assignment (COLA) metric [barrios2015cardinalized]. The main difference is that the COLA is not normalized by the cardinality. Another unnormalized OSPA-based metric is a variant of the COLA metric, misnamed as the Generalized OSPA (GOSPA) metric (which does not generalize OSPA in any way) [rahmathullah2017generalized]. Indeed, it is equivalent to the transport-transform (TT) metric proposed in [muller2020metrics], and thus, will be referred to as the TT metric since this is a more appropriate name. Apart from the mathematical similarity, it has been shown that COLA and TT metrics behave similarly in most scenarios, but TT is more sensitive to variations in parameters [barrios2023comparison].
@@
In many applications, each element of the point pattern represents a probability distribution on $\mathbb{R}^{n}$ to describe the associated uncertainty/confidence. Consequently, the discussed point pattern metrics need to be extended accordingly. Recall that the constructions of these metrics require a base metric or base distance between the elements. If the elements belong to $\mathbb{R}^{n}$, the $p$-norm of their difference is typically used as the base metric. When the elements are probability distributions, metrics such as Wasserstein [assa2018wasserstein], Hellinger, or geodesic distances on Gaussian manifolds [tang2018information, nielsen2023fisher] can be used as the base metric. Indeed, the Hellinger metric was used as a base metric in [nagappa2011incorporating] to incorporate estimation uncertainty into the OSPA metric. Further, information theoretic measures such as Kullback-Leibler’s divergence (KLD) [battistelli2014kullback], Jeffrey’s divergence (i.e., symmetrized KLD) [nielsen2009sided], Bhattacharyya distance [tang2021multi] can also be used to measure the dissimilarity between distributions, although they might not be proper metrics.
@@
In multi-object estimation, distances between multi-object trajectories are equally as important as distances between multi-object states. In [ristic2011metric] and [vu2014new], extensions of the OSPA metric (from multi-object states) to multi-object trajectories were proposed. However, these metrics are intractable, or are no longer proper metrics when numerically approximated in practice. A more practical approach is to treat a multi-object trajectory as a point pattern of trajectories, wherein point pattern metrics can be applied with tractable base distances between trajectories. A number of trajectory metrics were proposed in [beard2017ospa], but they exhibit counter-intuitive behavior, except for the OSPA-based construction. The OSPA metric with the OSPA-based trajectory base metric, called the OSPA(2) metric, addresses the shortcomings of previous multi-object trajectory metrics, offering meaningful physical interpretation and can be efficiently computed for large-scale scenarios [beard2020solution]. When the trajectories are sequences of probability distributions to describe the associated uncertainty, the OSPA-based trajectory metric is constructed from distances between probability distributions similar to that for the OSPA metric for point patterns of distributions discussed above. Further details and discussions on metrics for multi-object states and multi-object trajectories can be found in [nguyen2022trustworthy].
@@ III Trajectory and Fréchet Mean
This section presents a formulation and an accompanying numerical solution for trajectory consensus. Subsection III-A introduces the OSPA trajectory distance and formulates the Fréchet mean trajectory with respect to (w.r.t.) this distance. A method to compute the Fréchet mean trajectory based on greedy search is presented in Subsection III-B. A list of frequently used notation is given in Table I.
@@
$d_{\mathbb{X}}^{(c)}(\cdot,\cdot)$ a base distance on $\mathbb{X}$ cut-off at $c$
$d_{\mathbb{T}}^{(c)}(\cdot,\cdot)$ a trajectory distance on $\mathbb{T}$ cut-off at $c$
@@ III-A The Fréchet Mean Trajectory
Let $d_{\mathbb{X}}$ be a metric on the state space $\mathbb{X}$, and $d_{\mathbb{X}}^{(c)}(y,z)\triangleq\min\{d_{\mathbb{X}}(y,z),c\}$, $c>0$. For any trajectories $u$ and $v\in\mathbb{T}$, the OSPA trajectory distance with cut-off $c$ is defined as [beard2020solution]
$\!\!\!\!\!d_{\mathbb{T}}^{(c)}(u,v)=\\ \frac{{\scriptstyle{\displaystyle\sum_{k\in\mathcal{D}_{\!u}\cap\mathcal{D}_{\!v}}}\!}\!\!\!d_{\mathbb{X}}^{(c)}(u(k),v(k))+c(|\mathcal{D}_{\!u}\mathcal{-D}_{\!v}|+|\mathcal{D}_{\!v}\mathcal{-D}_{\!u}|)}{|\mathcal{D}_{\!u}\cup\mathcal{D}_{\!v}|}.\!\!$ (2)
This distance is well-defined because for any $u$, $v\in\mathbb{T}$, $|\mathcal{D}_{\!u}\cup\mathcal{D}_{\!v}|>0$, and satisfies all the metric properties [beard2020solution]. When the states of both trajectories exist (i.e., $k\in\mathcal{D}_{\!u}\cap\mathcal{D}_{\!v}$) the error is simply the distance between the respective states capped by $c$, and when only one state exists the error is $c$. Thus, intuitively the OSPA trajectory distance is simply the time-averaged state and cardinality errors between two trajectories.
With respect to the metric $d_{\mathbb{T}}^{(c)}(\cdot,\cdot)$, the $r^{th}$-order Fréchet mean of the trajectories $v^{(1)},...,v^{(N)}\in\mathbb{T}$ (hereon abbreviated as $v^{(1:N)}$) is
@@
[…]
@@
Suppose $\hat{v}$ is an $r^{th}$-order Fréchet mean of the trajectories $v^{(1:N)}$. Then for each $k\in\mathcal{D}_{\hat{v}}$, $\hat{v}(k)$ is the $r^{th}$-order Fréchet mean of the trajectory states $v^{(n)}(k)$ that exist at time $k$, weighted by $|\mathcal{D}_{\hat{v}}\cup\mathcal{D}_{v^{(n)}}|^{r}$, i.e.,
$\hat{v}(k)$ $=$ $\arg\min_{\mu\in\mathbb{X}}{\displaystyle\sum_{n:\mathcal{D}_{\!v^{(n)}\!}\ni k}}\frac{[d_{\mathbb{X}}^{(c)\!}(\mu,v^{(n)\!}(k))]^{r}}{|\mathcal{D}_{\hat{v}}\cup\mathcal{D}_{v^{(n)}}|^{r}}.$ (7)
The above result (see Supplementary Materials (SM) Subsection I-A for proof) provides an explicit expression for the state of the mean trajectory at each time step, given its domain. In the next subsection, this result will be used to compute the mean trajectory.

### v2

@@ Abstract
This paper introduces the concept of a mean for trajectories and multi-object trajectories (defined as sets or multi-sets of trajectories) along with algorithms for computing them. Specifically, we use the Fréchet mean, and metrics based on the optimal sub-pattern assignment (OSPA) construct, to extend the notion of average from vectors to trajectories and multi-object trajectories. Further, we develop efficient algorithms to compute these means using greedy search and Gibbs sampling. Using distributed multi-object tracking as an application, we demonstrate that the Fréchet mean approach to multi-object trajectory consensus significantly outperforms state-of-the-art distributed multi-object tracking methods.
@@ I Introduction
Multi-object estimation involves determining the trajectories of an unknown, time-varying number of objects from noisy sensor measurements. Formally, a trajectory is a function that maps time indices from a discrete time window to elements in some state space, and a multi-object trajectory is a point pattern or a multi-set [^1] A multi-set may contain repeated elements. of trajectories [beard2020solution]. This general definition covers trajectories that are fragmented (i.e., objects can disappear and then reappear later), which is usually the case with the output of most tracking systems. Estimating the multi-object trajectory lies at the heart of many applications ranging from space exploration [persico2018cubesat], surveillance [liu2022multi, nguyen2019trackerbots], computer vision [ma2024track, zhang2021fairmot], to cell biology [mavska2023cell, nguyen2021tracking].
@@
While it is customary that the mean is simply the sum of the data points divided by the number of data points, this is not the case in multi-object estimation. When the data points are trajectories of different lengths (see e.g., Figure 1), or multi-object trajectories (see e.g., Figure 2), the customary notion of average is not applicable because addition is not defined. The fundamental question then arises: what constitutes an “average” trajectory or an “average” multi-object trajectory? This question is crucial in multi-object estimation, yet, to the best of our knowledge, there is no existing work addressing the concept of an “average” trajectory or an “average” multi-object trajectory.
Figure 1: Two examples of the mean trajectory (thick black) of multiple input trajectories (color). Marker sizes indicate times, with the same size representing the same time, and sizes decrease toward the ends of trajectories. Consecutive instances of a trajectory are connected with solid lines.
Figure 2: Examples of the mean multi-object trajectory (MOT) of three sample multi-object trajectories. Each trajectory in a multi-object trajectory has a distinct color. Consecutive instances of a trajectory are connected with solid lines. Intuitively, the mean tends to smooth out the data points.
@@
Additionally, we demonstrate the capability of the proposed means in achieving consensus, using networked multi-object tracking as a specific area of application. Due to the proliferation of low-cost sensors that enable the deployment of large networks capable of monitoring large spatial regions, networked multi-object tracking is an emerging trend [iyer2022wind]. This approach leverages advances in communication and sensing technologies to improve accuracy by fusing information from multiple nodes in a flexible and reliable manner, resilient to node failures [luo2006distributed]. Given the local multi-object trajectory descriptions (e.g., estimates or probability distributions) at each node of the network, fusion is used to determine a multi-object trajectory description that best agrees with the local nodes, which can be posed as a consensus on the local multi-object trajectories. The proposed consensus can fuse node outputs regardless of the types of local trackers, and shows significant performance improvements over current state-of-the-art techniques.
@@ II-A Distributed Multi-Object Estimation
An alternative to multi-scan multi-object density fusion is track-to-track (T2T) fusion [chang1997optimal]. This is an intuitive technique for fusing individual tracks together by minimizing some prescribed cost functions. Additional information on the state can be augmented into the cost function [osbome2011track]. A popular cost function is based on the association probability of the sets of estimated objects with the physical objects [bar1981track, ogle2019alternative]. Alternatively, some authors use cost functions based on the statistical distance/divergence between the distributions describing the states of the tracks and their associated confidence [tang2021multi, battistelli2014kullback, tang2018information], or the OSPA distance between the tracks [nguyen2021distributed]. T2T assignment can be cast as a multi-dimensional assignment problem, which is NP-hard for more than two dimensions. Nevertheless, sub-optimal solutions can be obtained using efficient clustering techniques, including hierarchical clustering [bereson2006efficient], DBSCAN [he2018multi], density-peak clustering [chen2025distributed], or stochastic optimization [wolf2023track]. In T2T fusion, clutter objects are typically not considered within the optimization process but are instead discarded based on existence probability [muvsicki2015track], voting logic [zhang1997efficient], or assignment confidence [stankovic2021adaptive]. Determining an effective cost function for T2T fusion is still an active research topic [wang2022consistent, kaplan2008assignment, bar1986effect].
Numerically, multi-object trajectory fusion is more efficient with T2T fusion since only the individual states or their distributions are needed, rather than the entire multi-object densities. In [bar2004multisensor], the multi-scan assignment problem was solved using Lagrangian relaxation for a small number of sensors. A fast algorithm for multi-scan T2T assignment was proposed in [tokta2018fast], but it is limited to two sensors. The OSPA-based multi-object trajectory fusion approach proposed in [nguyen2021distributed] has been shown to be effective for a large number of sensors. Further, as a peer-to-peer fusion algorithm, it is more robust to network disruption. However, it is less accurate than global fusion algorithms, wherein all information is fused at once.
@@ II-B Metrics for Point Patterns and Trajectories
Since a multi-object state is a point pattern, metrics for point patterns are natural for measuring the distance between multi-object states. While the Hausdorff metric is an obvious point pattern metric, it is insensitive to cardinality differences [hoffman2004multitarget]. To address this, the optimal mass transfer metric was proposed in [hoffman2004multitarget], based on the Wasserstein construction. However, this metric suffers from a physical inconsistency if the cardinalities of the two multi-object states are not the same [hoffman2004multitarget]. A widely used multi-object metric with meaningful physical interpretations, capturing both the state and cardinality errors, is the OSPA metric [schuhmacher2008consistent]. The OSPA construction (by this we mean the use of optimal sub-pattern assignment as the abbreviation suggests) also admits other metrics such as the cardinalized optimal linear assignment (COLA) metric [barrios2017metrics]. The main difference is that the COLA is not normalized by the cardinality. Another unnormalized OSPA-based metric is a variant of the COLA metric, misnamed as the Generalized OSPA (GOSPA) metric [rahmathullah2017generalized], which does not generalize the OSPA metric [^2] GOSPA does not subsume the OSPA metric as a special case. In fact GOSPA is still based on the OSPA construct. . Indeed, it is equivalent to the transport-transform (TT) metric proposed in [muller2020metrics], and thus, will be referred to as the TT metric since this is a more appropriate name. Apart from the mathematical similarity, it has been shown that COLA and TT metrics behave similarly in most scenarios, but TT is more sensitive to variations in parameters [barrios2023comparison].
The TT and OSPA metrics are natural candidate distances to formulate the Fréchet mean of point patterns. Techniques for computing point pattern barycentres w.r.t. the TT and OSPA metrics have been discussed in [muller2020metrics]. An earlier approach for computing the OSPA barycenter using an alternating greedy search algorithm was proposed in [baum2015ospa]. This technique can be applied to approximate the minimum mean OSPA estimate (from certain distributions), first formulated in [guerriero2010shooting] for fixed cardinality, and later extended to unknown cardinality in [balasingam2015mmospa].
@@
In many applications, each element of the point pattern represents a probability distribution on $\mathbb{R}^{n}$ to describe the associated uncertainty/confidence. Consequently, the discussed point pattern metrics need to be extended accordingly. Recall that the constructions of these metrics require a base metric or base distance between the elements. If the elements belong to $\mathbb{R}^{n}$, the $p$-norm of their difference is typically used as the base metric. When the elements are probability distributions, metrics such as Wasserstein [assa2018wasserstein], Hellinger, or geodesic distances on Gaussian manifolds [tang2018information, nielsen2023fisher] can be used as the base metric. Indeed, the Hellinger metric was used as a base metric in [nagappa2011incorporating] to incorporate estimation uncertainty into the OSPA metric. Further, information theoretic measures such as Kullback-Leibler’s divergence (KLD) [battistelli2014kullback], Jeffrey’s divergence (i.e., symmetrized KLD) [nielsen2009sided], Bhattacharyya distance [tang2021multi] can also be used to measure the dissimilarity between distributions, though they may not satisfy all metric properties.
@@
In multi-object estimation, distances between multi-object trajectories are equally as important as distances between multi-object states. In [ristic2011metric] and [vu2014new], extensions of the OSPA metric (from multi-object states) to multi-object trajectories were proposed. However, these metrics are intractable, or are no longer proper metrics when numerically approximated in practice. A more practical approach is to treat a multi-object trajectory as a point pattern of trajectories, wherein point pattern metrics can be applied with tractable base distances between trajectories. A number of trajectory metrics were proposed in [beard2017ospa], but they exhibit counter-intuitive behavior, except for the OSPA-based construction. The OSPA metric with the OSPA-based trajectory base metric, called the OSPA(2) metric, addresses the shortcomings of previous multi-object trajectory metrics, offering meaningful physical interpretation and can be efficiently computed for large-scale scenarios [beard2020solution]. The TGOSPA metric [garcia2020ametric], proposed after OSPA(2), is not applicable in this work, because it was only proven for the special case of multi-object trajectories with contiguous trajectories. When the trajectories are sequences of probability distributions to describe the associated uncertainty, the OSPA-based trajectory metric is constructed from distances between probability distributions similar to that for the OSPA metric for point patterns of distributions discussed above. Further details and discussions on metrics for multi-object states and multi-object trajectories can be found in [nguyen2022trustworthy].
@@ III Trajectory and Fréchet Mean
This section presents a formulation and an accompanying numerical solution for trajectory consensus. Subsection III-A introduces the OSPA trajectory distance and formulates the Fréchet mean trajectory w.r.t. this distance. A method to compute the Fréchet mean trajectory based on greedy search is presented in Subsection III-B. A list of frequently used notation is given in Table I.
@@
$d_{\mathbb{X}}^{(c)}(\cdot,\cdot)$ a base distance on $\mathbb{X}$ with cut-off at $c$
$d_{\mathbb{T}}^{(c)}(\cdot,\cdot)$ a trajectory distance on $\mathbb{T}$ with cut-off at $c$
@@ III-A The Fréchet Mean Trajectory
Let $d_{\mathbb{X}}$ be a metric on the state space $\mathbb{X}$, and $d_{\mathbb{X}}^{(c)}(y,z)\triangleq\min\{d_{\mathbb{X}}(y,z),c\}$, $c>0$. For any trajectories $u$ and $v\in\mathbb{T}$, the OSPA trajectory distance of order $r\in[1,\infty]$ and cut-off $c$ is defined as
$d_{\mathbb{T}}^{(c,r)}(u,v)=\\ \!\!\!\left[\!\frac{{\scriptstyle{\displaystyle\sum_{k\in\mathcal{D}_{\!u}\cap\mathcal{D}_{\!v}}}\!}\!\!\![d_{\mathbb{X}}^{(c)}(u(k),v(k))]^{r}\!+\!c^{r}(|\mathcal{D}_{\!u}\mathcal{-D}_{\!v}|+|\mathcal{D}_{\!v}\mathcal{-D}_{\!u}|)}{|\mathcal{D}_{\!u}\cup\mathcal{D}_{\!v}|}\!\right]^{\!\frac{1}{r}}\!\!.\!\!$ (2)
Remark 1.
The proof of the OSPA trajectory metric for $r=1$ was given in [beard2020solution]. Nonetheless, this result holds for any $r\in[1,\infty]$, and for completeness see Supplementary Materials (SM) Subsection I-A for proof. When the states of both trajectories exist (i.e., $k\in\mathcal{D}_{\!u}\cap\mathcal{D}_{\!v}$) the error is simply the distance between the respective states capped by $c$, and when only one state exists the error is $c$. Thus, intuitively the OSPA trajectory distance is simply the $r^{th}$-order time-averaged state and cardinality errors between two trajectories.
With respect to the metric $d_{\mathbb{T}}^{(c,r)}(\cdot,\cdot)$, the $r^{th}$-order Fréchet mean of the trajectories $v^{(1)},...,v^{(N)}\in\mathbb{T}$ (hereon abbreviated as $v^{(1:N)}$) is
@@
$V^{(r)}(u)$ $\!\!\!\!=\!\!\!\!$ ${\displaystyle\sum_{k\in\mathbb{K}}}\mathbf{1}_{\mathcal{D}_{u}}^{(k)}\Psi_{k}^{(r)}(u)+\mathbf{1}_{\mathcal{\overline{D}}_{u}}^{(k)}\bar{\Psi}_{k}^{(r)}(u),$ (4)
$\Psi_{k}^{(r)}(u)$ $\!\!\!\!=\!\!\!\!$ ${\displaystyle\sum_{n=1}^{N}}\frac{\mathbf{1}_{\mathcal{D}_{\!v^{(n)}\!}}^{(k)}[d_{\mathbb{X}}^{(c)\!}(u(k),v^{(n)\!}(k))]^{r}+\mathbf{1}_{\mathcal{\overline{D}}_{\!v^{(n)}\!}}^{(k)}c^{r}}{|\mathcal{D}_{u}\cup\mathcal{D}_{v^{(n)}}|},$ (5)
$\bar{\Psi}_{k}^{(r)}(u)$ $\!\!\!\!=\!\!\!\!$ ${\displaystyle\sum_{n=1}^{N}}\frac{\mathbf{1}_{\mathcal{D}_{\!v^{(n)}}}^{(k)}c^{r}}{|\mathcal{D}_{u}\cup\mathcal{D}_{\!v^{(n)}}|},$ (6)
@@
Suppose $\hat{v}$ is an $r^{th}$-order Fréchet mean of the trajectories $v^{(1:N)}$. Then for each $k\in\mathcal{D}_{\hat{v}}$, $\hat{v}(k)$ is the $r^{th}$-order Fréchet mean of the trajectory states $v^{(n)}(k)$ that exist at time $k$, weighted by $|\mathcal{D}_{\hat{v}}\cup\mathcal{D}_{v^{(n)}}|$, i.e.,
$\hat{v}(k)$ $=$ $\arg\min_{\mu\in\mathbb{X}}{\displaystyle\sum_{n:\mathcal{D}_{\!v^{(n)}\!}\ni k}}\frac{[d_{\mathbb{X}}^{(c)\!}(\mu,v^{(n)\!}(k))]^{r}}{|\mathcal{D}_{\hat{v}}\cup\mathcal{D}_{v^{(n)}}|}.$ (7)
The above result (see SM Subsection I-B for proof) provides an explicit expression for the state of the mean trajectory at each time step, given its domain. In the next subsection, this result will be used to compute the mean trajectory.